Accounts receivable turnover

10:1

Total assets turnover

2:1

Average receivables during the year

₱400,000

Total assets, January 1, 20x1

800,000


How much is the total assets as of December 31, 20x1?

a. 4,000,000
b. 3,800,000
c. 3,200,000
d. 2,800,000
Answer #1

Accounts receivable turnover = Sales / Account receivable

10 = Sales / 400000

Sales = 10 * 400000

= 4000000

Total assets turnover = Sales / Average assets

2 = 4000000 / Average assets

Average assets = 4000000 / 2

= 2000000

Average assets = (Beginning assets + Ending assets) / 2

2000000 = (800000 + Ending assets) / 2

Ending assets = 2000000 * 2 - 800000

= 3200000

Correct choice C